# Post-detection inference for sequential changepoint localization

### v1

@@ A.2 Omitted proofs from Section 3
Define, $C_{t,L}^{\alpha}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\alpha;M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$. Note that $\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1}$ imply that $C_{t,L}^{\alpha}\geq C_{t,\infty}^{\alpha}\geq\operatorname{Quantile}(1-\alpha;M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B}\}),$ for all $t\in\mathbb{N},\alpha\in(0,1)$. Also, for all $t$, with $r_{t}=\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$, since $r_{t}$ is independent to $\{X_{n}\}_{n}$, we have
$\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})\mid r_{t}))\geq\mathbb{E}(1-\beta r_{t})=1-\beta\mathbb{P}_{\theta_{0},\infty}(\tau\geq t).$
@@
$\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B}))+\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
$\leq\mathbb{E}\left(\mathbb{P}_{\theta_{0},t,\theta_{1}}\left(M_{t}>\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B})\mid\hat{\alpha}_{t}\right)\right)+\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
$\leq\mathbb{E}(\hat{\alpha}_{t})+\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
$=\alpha\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)+\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)\leq(\alpha+\beta)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t).$
Similarly, we have for fixed $\alpha\in(0,1),$ with $\mathcal{S}^{\prime}_{\tau-t+1}$ having coverage $1-\beta$
$P_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha})\leq\alpha+\beta.$ (A.4)
@@
$=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}}\}\mid\tau\geq T)\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau\geq T)+\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}}\},\tau<T)$
Similarly, we have
$P_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha})\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+1-\mathbb{P}_{\theta_{0},\infty}(\tau\geq T).$ (A.5)
@@ A.3 Omitted proofs from Section 4
Define, $\hat{\alpha}_{t}:=\frac{\alpha}{N}\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)$ and $C_{t,L}^{\alpha}=\operatorname{Quantile}(1-\alpha;M_{t},\{\max_{\theta\in\mathcal{S}_{t-1},\theta^{\prime}\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$. Note that $\theta_{0}\in\mathcal{S}_{t-1}$ and $\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1}$ imply that $C_{t,L}^{\alpha}\geq C_{t,\infty}^{\alpha}\geq\operatorname{Quantile}(1-\alpha;M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}\}),$ for any $\alpha\in(0,1).$ By Assumption 4.1, we have that $\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)\geq\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$.
Also, for all $t$, with $r_{t}=\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$, since $r_{t}$ is independent to $\{X_{n}\}_{n}$, we have
$\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})\mid r_{t}))\geq\mathbb{E}(1-\beta r_{t})=1-\beta\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t).$
@@
$\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{0}\in\mathcal{S}_{t-1})=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{0}\in\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t})\mid r_{t}))\geq\mathbb{E}(1-\gamma r_{t})=1-\gamma\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t).$
@@
$\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\hat{\alpha}_{t}})$
$\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\hat{\alpha}_{t}},\theta_{0}\in\mathcal{S}_{t-1},\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})+\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{0}\notin\mathcal{S}_{t-1}))+\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\notin\mathcal{S}^{\prime}_{\tau-t+1})$
$\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}))+(\beta+\gamma)\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$
$\leq\mathbb{E}\left(\mathbb{P}_{\theta_{0},t,\theta_{1}}\left(M_{t}>\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B})\mid\hat{\alpha}_{t}\right)\right)+(\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
$\leq\mathbb{E}(\hat{\alpha}_{t})+(\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
@@
Similarly, we have for fixed $\alpha\in(0,1),$ with $\mathcal{S}_{t-1}$ and $\mathcal{S}^{\prime}_{\tau-t+1}$ having coverages $1-\beta$ and $1-\gamma$ respectively,
$P_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha})\leq\alpha+\beta+\gamma.$ (A.7)
@@
$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\hat{\alpha}_{t}}\},\tau<T)\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau<T),$
@@
$\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{t,L}^{\hat{\alpha}_{t}})$
$=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\hat{\alpha}_{t}}\})$
$=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\hat{\alpha}_{t}}\}\mid\tau\geq T)\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau\geq T)+\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\hat{\alpha}_{t}}\},\tau<T)$
@@
Since $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{t,L})\geq 1-(\alpha+\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)$, we obtain $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta-\gamma$ from the above inequality. Similarly, we have
$P_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha})\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+1-\mathbb{P}_{\theta_{0},\infty}(\tau\geq T).$ (A.8)

### v2

@@ A.2 Omitted proofs from Section 3
Define, $C_{t,L}^{\alpha}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\alpha;M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$. Note that $\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1}$ imply that $C_{t,L}^{\alpha}\geq C_{t,\infty}^{\alpha}\geq\operatorname{Quantile}(1-\alpha;M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B}\}),$ for all $t\in\mathbb{N},\alpha\in(0,1)$.
  1. (a)
    For $\mathcal{C}_{1}$: For all $t$, with $r_{t}=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$, since $r_{t}$ is independent to $\{X_{n}\}_{n}$, we have
    $\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})$ $=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})\mid r_{t}))$
    $\geq\mathbb{E}(1-\beta r_{t})=1-\beta\mathbb{P}_{\theta_{0},\infty}(\tau\geq t).$
@@
    $\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B}))+\beta\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
    $\leq\mathbb{E}\left(\mathbb{P}_{\theta_{0},t,\theta_{1}}\left(M_{t}>\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B})\mid r_{t}\right)\right)+\beta\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
    $\leq\mathbb{E}(\alpha r_{t})+\beta\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
    $=(\alpha+\beta)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t).$
@@
    $\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}}\}\mid\tau\geq T)\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau\geq T)+\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau<T)$
    $=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+1-\mathbb{P}_{\theta_{0},\infty}(\tau\geq T).$
@@
  2. (b)
    For $\mathcal{C}_{2}$: Here, we have $r_{t}=1$. The proof for $\mathcal{C}_{2}$ is similar. For fixed $\alpha\in(0,1),$ with $\mathcal{S}^{\prime}_{\tau-t+1}$ having coverage $1-\beta$ we show that
    $P_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha})\leq\alpha+\beta.$
    Then, we have
    $P_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha})\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+1-\mathbb{P}_{\theta_{0},\infty}(\tau\geq T).$
    If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, we obtain
@@ A.3 Omitted proofs from Section 4
Define, $C_{t,L}^{\alpha}=\operatorname{Quantile}(1-\alpha;M_{t},\{\max_{\theta\in\mathcal{S}_{t-1},\theta^{\prime}\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$. Note that $\theta_{0}\in\mathcal{S}_{t-1}$ and $\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1}$ imply that $C_{t,L}^{\alpha}\geq C_{t,\infty}^{\alpha}\geq\operatorname{Quantile}(1-\alpha;M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}\}),$ for any $\alpha\in(0,1).$ By Assumption 4.1, we have that $\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)\geq\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$.
  1. (a)
    For $\mathcal{C}_{1}$: Also, for all $t$, with $r_{t}=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$, since $r_{t}$ is independent to $\{X_{n}\}_{n}$, we have
    $\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})$ $=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})\mid r_{t}))$
    $\geq\mathbb{E}(1-\beta r_{t})=1-\beta\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t).$
@@
    $\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{o}\in\mathcal{S}_{t-1})$ $=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t})\mid r_{t}))$
    $\geq\mathbb{E}(1-\gamma r_{t})=1-\gamma\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t).$
@@
    $\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha r_{t}})$
    $\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha r_{t}},\theta_{0}\in\mathcal{S}_{t-1},\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})+\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{0}\notin\mathcal{S}_{t-1}))+\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\notin\mathcal{S}^{\prime}_{\tau-t+1})$
    $\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}))+(\beta+\gamma)\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$
    $\leq\mathbb{E}\left(\mathbb{P}_{\theta_{0},t,\theta_{1}}\left(M_{t}>\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B})\mid r_{t}\right)\right)+(\beta+\gamma)\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
    $\leq\mathbb{E}(\alpha r_{t})+(\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
@@
    $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}}\},\tau<T)\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau<T),$
@@
    $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{t,L}^{\alpha r_{t}})$
    $=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}}\}\mid\tau\geq T)\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau\geq T)+\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau<T)$
@@
    Since $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{t,L})\geq 1-(\alpha+\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)$, we obtain $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta-\gamma$ from the above inequality.
  2. (b)
    For $\mathcal{C}_{2}$: Proof is similar to part (a). Here we have for fixed $r_{t}=1$ with $\mathcal{S}_{t-1}$ and $\mathcal{S}^{\prime}_{\tau-t+1}$ having coverages $1-\beta$ and $1-\gamma$ respectively,
    $P_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha})\leq\alpha+\beta+\gamma.$
    Similarly, we have
    $P_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha})\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+1-\mathbb{P}_{\theta_{0},\infty}(\tau\geq T).$